Fe2O3 -> Fe3O4 ,In this reaction what is the equivalent weight of Fe2O3 ?

Grade:12th pass

1 Answers

Arun
4 years ago

The oxidation state of Fe in Fe²O³ is +3.

3 – 8/3 = 1/3

That is 1/3 mole of electron is required

Therefore, the equivalent weight of iron in Fe²O³ = 55.8 / (1/3) = 167.4 gm